Find the vertical displacement of joint C. Assume a= 1.5 m, b = 2 m, P= 5 kN. AE = const for all bars. B Р
Answer #1

O= 45° Ay + Dy =P Ay=Dy = 2 (1) Dx=0 at joint à fy=0 - I - FAC Cassin4520 & FAB / FAC = f (T) {&x=0 & FAR & FAC Sinus =0 ragavertically apply unit load at plc in downwards This is similar to previous 212 if p=1N c face too 1.5 212 I 1-5 PAR =- SL FBB